OpenStudy (rose4825):

What is the taxidistance between (2, 11) and (7, 10)? 7 4 6 26 My answer: 6

OpenStudy (rose4825):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x difference: 7-2 = 5 y difference: 11-10 = 1 Add up those results: 5+1 = 6 again you have the correct answer